# Association and Independence Test for Random Objects

### v2

@@ 6.1 Connectivity networks in fMRI based brain imaging
The data are publicly accessible at https://www.humanconnectome.org. For our analysis, we focus on a subset comprising $n=209$ young, healthy adults who completed both resting-state and motor task fMRI sessions. The resting-state fMRI data were collected while participants were instructed to keep their eyes open and fixate on a projected bright cross-hair on a dark background in a darkened room. For the motor task fMRI recordings, participants were presented with visual cues prompting them to perform specific movements, such as squeezing their left or right toes or moving their tongue, to map specific motor areas. Each time block for the recording of a specific movement type lasted 12 seconds, included 10 movements and was preceded by a 3-second cue. After preprocessing, the average blood-oxygen-level-dependent (BOLD) signals were recorded across different voxels at 284 time points. Further details about the experiment, preprocessing and data acquisition are available in the reference manual for the WU-Minn HCP 1200 Subjects Data Release (Wu-Minn, 2017).
@@
Our analysis focuses on nine regions of interest, nucleus accumbens, amygdala, caudate, cerebellum, diencephalon, hippocampus, pallidum, putamen, and thalamus, in both left and right brain hemispheres. Correlation matrices of BOLD signals were constructed as per established practice in fMRI analysis, namely calculating the normalized cross-correlations of fMRI time series based on the voxels residing around a seed region for each region of interest, separately for left and right brain hemispheres. For the task-evoked fMRI signals, the correlation matrices were calculated using the time segments corresponding to each specific task. For instance, the correlation matrix for “squeezing left toe” was derived from the time periods during which participants were actively performing the “squeezing left toe” task. For each subject, this resulted in a total of eight $9\times 9$ correlation matrices, corresponding to one rest phase and three task phases (resting state, squeezing left toe, squeezing right toe, and moving tongue) for each of the two brain hemispheres (left and right).
Figure 3 shows the estimated Fréchet mean of these eight covariance matrices when employing the Frobenius norm. This indicates that the task-evoked correlation matrices display higher correlations between different brain regions compared with the resting-state correlation matrices.
@@
Figure 3: Fréchet mean correlation matrices for resting-state fMRI (left-most column) and task-evoked fMRI (next three columns: squeezing left toe, squeezing right toe, and moving tongue). The top row represents the left hemisphere and the bottom row the right hemisphere.
@@
We then proceeded to apply the proposed profile independence test, along with the energy test (Székely et al., 2007), aiming to test the independence of random object pairs $\{(X_{i},Y_{i})\}_{i=1}^{209}$. Here $X_{i}$ represents the correlation matrix corresponding to resting-state fMRI, and $Y_{i}$ the correlation matrix corresponding to task-evoked fMRI, for the various motor tasks for the $i$th subject. We consider three metrics for the space of $9\times 9$ positive definite matrices $A$ and $B$. Specifically, we compare the results for the following three metrics that were applied previously in the literature for various data analysis tasks:
@@
For each metric and method, we conducted six tests in total: three for each motor task (“squeezing the left toe”, “squeezing the right toe”, and “moving the tongue”) separately for the left and right hemispheres. The $p$-values resulting from the permutation test implementation of the proposed profile independence test are shown in Table 1 and are also compared with the results from the energy test for independence. For the tongue movement task (columns 3, 6, and 9), both the proposed profile independence test and the energy test yield the same result across all three metrics: there is no evidence to reject the null hypothesis that the correlation matrices between the resting state and the tongue movement task are independent. For the toe-squeezing task, under the Frobenius metric (columns 1 and 2), both the proposed profile independence test and the energy test indicate dependence at the 0.05 significance level. However, the energy test fails to detect dependence under the AIRM and Log-Cholesky metrics (row 2, columns 5 and 8), whereas the proposed profile independence test consistently rejects the null hypothesis in all these cases. These results demonstrate the consistency of the proposed test across different metric choices.
Table 1: Comparison of $p$-values obtained when applying the proposed profile independence test and the energy test for independence for the null hypothesis of independence between resting-state fMRI and task-evoked fMRI connectivity matrices for three motor tasks: $\mathrm{Ltoe}$ (squeezing the left toe), $\mathrm{Rtoe}$ (squeezing the right toe), and $\mathrm{tongue}$ (moving the tongue). The results are presented for left and right hemispheres and for three different metrics: Frobenius metric (left three columns), Affine-invariant Riemannian metric (middle three columns) and Log-Cholesky metric (right three columns). The first two rows correspond to correlation matrices show the results for the left hemisphere and the last two rows for the right hemisphere. The $p$-values for the proposed profile independence test and the energy test were obtained through permutation tests with a total number of permutations $N=500$ and are bolded if $p>0.1$.
@@ 6.2 Age-at-death distributions
The study of human mortality as quantified by age-at-death distributions has attracted much interest over the past decades. The analysis of age-at-death distributions has provided insights into mortality patterns, lifespan variability and the factors influencing survival across different populations and time periods. The Human Mortality Database (www.mortality.org) provides annual age-at-death tables for 38 countries, with death rates recorded for each country across ages ranging from 0 to 110. By applying basic smoothing techniques to the yearly life tables available in this database, one may obtain smooth age-at-death density distributions indexed by both country and calendar year. We focus on 33 countries with available data spanning the years 1983 to 2018 and employ the conditional framework detailed in Section 4, where the data are considered to be triplets $(X,Y,Z)$. Here $X$ stands for the male age-at-death distribution for a given country at calendar year $Z$, and $Y$ the corresponding age-at-death distribution for females for the same country and year. We employ the Wasserstein-2 metric (16) in the space of one-dimensional distributions considered here, as it is particularly sensitive to horizontal shifts in distributions that are of special interest.
@@
We apply the proposed conditional profile association $\Delta_{XY}(z)$ in (12) for this analysis, aiming to shed light on the changing nature of the relationship between male and female age-at-death distributions across different calendar years. The conditional profile association serves to quantify the degree of dependence between the two distributions as a function of calendar time $z$, where a smaller value of $\Delta_{XY}(z)$ indicate that male and female death rate distributions are relatively less dependent. dependent, while larger values indicate stronger association and a higher degree of dependence.
@@
We divide the 33 countries into two groups: Eastern European countries, which include Bulgaria, Belarus, Czechia, Estonia, Hungary, Lithuania, Latvia, Poland, Slovakia, and Slovenia; and other countries that can be broadly characterized as Western/Asian and include Australia, Austria, Belgium, Canada, Switzerland, Denmark, Spain, Finland, France, England and Wales, Greece, Ireland, Iceland, Italy, Japan, Luxembourg, the Netherlands, Norway, New Zealand, Portugal, Sweden, Taiwan and the United States. The results are in Figure 4, which illustrates the estimated conditional profile association as a function of calendar year for all countries combined, as well as separately for Eastern European and Western/Asian countries.
@@
The left panel, representing all countries, exhibits a generally decreasing trend over time, indicating a gradual reduction in the association between male and female age-at-death distributions over the 25 years covered in this analysis. The middle panel, which focuses on Eastern European countries only, shows a noticeable peak in the early 1990s, suggesting a period of heightened dependence between male and female mortality distributions, potentially due to major socio-political and economic changes in the region during that time that had a strong effect on both female and male age-at-death distributions. After this peak, the association declines but remains relatively stable. In contrast, the right panel, which represents Western/Asian countries, shows a consistently lower conditional profile association compared to both all countries combined and Eastern European countries. The association in Western countries also follows a steady downward trend, suggesting an increasing independence between male and female mortality distributions over time. This lower and declining association seems to indicate an increasing decoupling between male and female mortality over time that is much more expressed for Western/Asian countries but is much less evident for Eastern European countries. The clear disparity between Western and Eastern European countries highlights the varying demographic, political and health-related forces that influence the relationship between male and female mortality.
@@ 7 Discussion
In this paper, we introduce the profile association, a measure designed to quantify associations between random objects. The proposed association is based on distance profiles, which serve as bridges connecting the distributions of random objects, the classical Hoeffding’s D statistic, and the proposed association measure. Utilizing this framework, we develop an independence testing procedure for random objects, demonstrating superior performance compared to existing methods in numerical studies. We further extend the profile association and the associated testing framework to conditional settings, in which random objects are paired with Euclidean predictors. This extension provides a useful tool for exploring relationships between complex object responses and Euclidean covariates, with immediate practical relevance. Beyond independence testing, the profile association framework has potential for broader statistical applications, including causal inference, change-point detection, and feature selection, thus opening promising avenues for future research.
Causal inference has emerged as an important field in statistics, with significant applications across various scientific disciplines. One key challenge in causal inference is detecting confounding contributions, which represent spurious associations arising from the influence of a confounder on two outcomes of interest. A simple conceptual tool for understanding confounding for Euclidean data is the classical law of total covariance,
@@
which decomposes the conditional covariance between random variables $X$ and $Y$ in the presence of a conditioning variable $Z$. Here $\text{Cov}(\mathbb{E}[X\mid Z],\mathbb{E}[Y\mid Z])$ is the contribution to the covariance explained by $Z$, and thus possibly due to a confounding effectof $Z$, whereas $\text{Cov}(\mathbb{E}[X\mid Z],\mathbb{E}[Y\mid Z])$ is the part that is not explained by $Z$ in the context of a linear framework.
An interesting question for future research is whether by utilizing the proposed conditional profile association one can extend such a separation of the effect that can be attributed to $Z$ and the remaining association that cannot be explained by $Z$ to scenarios where $X$ and $Y$ are random objects and a linear framework does not apply, due to the lack of vector operations in non-Euclidean spaces. Considering the mortality example discussed in Section 6.2, where $(X_{i}$ and $Y_{i}$ are female and male age-at-death distributions, while $Z_{i}$ is a Euclidean covariate, a preliminary attempt to shed some light on this is the following three-step procedure: first compute the unconditional profile association $\Delta(X,Y)$, using all available data pairs $\{(X_{i},Y_{i})\}_{i}$; second, replace the conditional expectations $\mathbb{E}[X\mid Z]$, $\mathbb{E}[Y\mid Z])$ in the law of total covariance by Fréchet regressions (Petersen and Müller, 2019) to obtain conditional Fréchet means $\mathbb{E}_{\oplus}[X\mid Z]$ and $\mathbb{E}_{\oplus}[Y\mid Z]$ using data $\{(X_{i},Z_{i})\}_{i}$ and $\{(Y_{i},Z_{i})\}_{i}$ respectively; third, compute $\Delta(\mathbb{E}_{\oplus}[X\mid Z],\mathbb{E}_{\oplus}[Y\mid Z])$ by plugging in the results from the Fréchet regression step.
Applying these considerations for the human mortality data for all countries combined yields $0.3630\Delta(X,Y)=$ 0.36 for the estimate of the overall profile association that ignores calendar year, but $30\Delta(\mathbb{E}[_{\oplus}X\mid Z],\mathbb{E}[_{\oplus}Y\mid Z])=0.97$ for the estimate of the part of the association explained by $Z$. Here we introduce the multiplier $30$ to normalize the association values so they lie in the interval $[0,1]$. Thus when calendar year is ignored, the observed profile association between male and female age-at-death distributions is much weaker than the value obtained when one first uses the calendar year as predictor. This discussion is not meant to be conclusive but rather to showcase the potential for future statistical research on theory and applications of conditional profile association for random objects.

### v3

@@ 6.1 Connectivity networks in fMRI based brain imaging
The data are publicly accessible at https://www.humanconnectome.org. For our analysis, we focus on a subset comprising $n=209$ young healthy adults who completed both resting-state and motor task fMRI sessions. The resting-state fMRI data were collected while participants were instructed to keep their eyes open and fixate on a projected bright cross-hair on a dark background. For the motor task fMRI recordings, participants were presented with visual cues prompting them to perform specific movements, such as squeezing their left or right toes or moving their tongue. Each time block for the recording of a specific movement type lasted 12 seconds, included 10 movements and was preceded by a 3-second cue. After preprocessing, the average blood-oxygen-level-dependent (BOLD) signals were recorded across different voxels at 284 time points. Further details about the experiment, preprocessing and data acquisition are available in the reference manual for the WU-Minn HCP 1200 Subjects Data Release (Wu-Minn, 2017).
@@
Our analysis focuses on nine regions of interest, nucleus accumbens, amygdala, caudate, cerebellum, diencephalon, hippocampus, pallidum, putamen, and thalamus, in both left and right brain hemispheres. Correlation matrices of BOLD signals were constructed as per established practice in fMRI analysis, namely calculating the normalized cross-correlations of fMRI time series based on the voxels residing around a seed region for each region of interest, separately for left and right brain hemispheres. For the task-evoked fMRI signals, the correlation matrices were calculated using the time segments corresponding to each specific task. For instance, the correlation matrix for “squeezing left toe” was derived from the time periods during which participants were actively performing the “squeezing left toe” task. For each subject, this resulted in a total of four $18\times 18$ correlation matrices for the whole brain, corresponding to one rest phase and three task phases (resting state, squeezing left toe, squeezing right toe, and moving tongue).
Figure 4 shows the estimated Fréchet mean of the four covariance matrices under the Frobenius norm. The results indicate that task-evoked correlation matrices exhibit stronger correlations across different brain regions compared to resting-state correlation matrices. Furthermore, in both resting-state and task-evoked conditions, the connectivity matrices between the left and right hemispheres are highly correlated. Applying both the proposed independence test and the energy test to these hemispheric connectivity matrices yields $p$-values below 0.01 in all cases, confirming that the left and right brain connectivities are not independent, which is consistent with expectations.
@@
Figure 4: Fréchet mean correlation matrices for resting-state fMRI (leftmost column) and task-evoked fMRI (next three columns: squeezing the left toe, squeezing the right toe, and moving the tongue). In each matrix, the top-left $9\times 9$ block represents the connectivity within the left hemisphere, and the bottom-right $9\times 9$ block represents the connectivity within the right hemisphere.
@@
Given the strong dependence between left and right hemisphere connectivity matrices observed across both resting-state and task-evoked conditions, we focus our analysis on separate $9\times 9$ intra-hemispheric correlation matrices. This is motivated by the fact that when left and right hemispheres are highly correlated, performing independence tests on the full $18\times 18$ whole-brain matrices can make it easier to reject the null hypothesis of independence due to inflated signal from inter-hemispheric correlations. In contrast, tests based on the smaller $9\times 9$ matrices are more stringent and informative for evaluating within-hemisphere dynamics. To this end, we applied the proposed test and the energy test (Székely et al., 2007) to assess the independence of random object pairs $\{(X_{i},Y_{i})\}_{i=1}^{209}$. Here, $X_{i}$ denotes the correlation matrix from the resting-state fMRI, and $Y_{i}$ denotes the correlation matrix from the task-evoked fMRI corresponding to the various motor tasks, for the left and right hemispheres of the $i$th subject. We consider three metrics for the space of $9\times 9$ positive definite matrices $A$ and $B$. Specifically, we compare the results for the following metrics that were applied previously in the literature for various data analysis tasks:
@@
  • c)
    $\alpha$-Power metric (Dryden et al., 2009; Zhou and Mülller, 2022)
    $d_{F,\alpha}=\|A^{\alpha}-B^{\alpha}\|_{F},$
    where $A^{\alpha}=U\Lambda^{\alpha}U^{\top}$ and $A=U\Lambda U^{\top}$ is the usual spectraL decomposition of a positive definite matrix AA.
  • d)
    Bures-Wasserstein metric
    $d_{{BW}}(A,B)^{2}=\operatorname{Tr}(A+B-2(A^{1/2}BA^{1/2})^{1/2}).$
For each metric and method, we conducted six tests in total: three for each motor task (“squeezing the left toe”, “squeezing the right toe”, and “moving the tongue”) separately for the left and right hemispheres. The $p$-values resulting from the permutation test implementation of the proposed profile independence test are shown in Table 1 and are also compared with the results from the energy test for independence. For the tongue movement task, both the proposed profile independence test and the energy test yield the same result across all three metrics: there is no evidence to reject the null hypothesis that the correlation matrices between the resting state and the tongue movement task are independent. For the toe-squeezing task, under the Frobenius metric, both the proposed profile independence test and the energy test indicate dependence at the 0.05 significance level. However, the energy test fails to detect dependence under the AIRM, Log-Cholesky, $1/2$-Power, $1/3$-Power and Bures–Wasserstein metrics for the right toe-squeezing task, whereas the proposed profile independence test consistently rejects the null hypothesis in all these cases. These results demonstrate the consistency of the proposed test across different metric choices.
For completeness, we also include test results based on the full $18\times 18$ correlation matrices in the Supplement, with the above caveat. Both the proposed method and the energy test yield consistently low $p$-values for the left and right toe-squeezing tasks, whereas the $p$-values for the tongue movement task are comparatively higher. These results are consistent across all metrics.
Table 1: Comparison of $p$-values obtained when applying the proposed profile independence test and the energy test for independence for the null hypothesis of independence between resting-state fMRI and task-evoked fMRI connectivity matrices for three motor tasks: $\mathrm{Ltoe}$ (squeezing the left toe), $\mathrm{Rtoe}$ (squeezing the right toe), and $\mathrm{tongue}$ (moving the tongue). The results are presented for left and right hemispheres and for five different metrics: Frobenius metric (left three columns, upper table), Affine-invariant Riemannian metric (middle three columns, upper table), Log-Cholesky metric (right three columns, upper table), $1/2$-Power metric (left three columns, lower table), $1/3$-Power metric (middle three columns, lower table) and Bures–Wasserstein metric (right three columns, lower table). The first two rows in upper/lower table show the results for the correlation matrices of the left hemisphere and the last two sub-rows for the right hemisphere. The $p$-values for the proposed profile independence test and the energy test were obtained through permutation tests with a total number of permutations $N=500$ and are bolded if $p>0.1$.
@@
1/2-Power metric 1/3-Power metric Bures-Wasserstein
Ltoe Rtoe tongue Ltoe Rtoe tongue Ltoe Rtoe tongue
Left proposed 0.00 0.03 0.55 0.00 0.02 0.68 0.00 0.03 0.54
energy 0.01 0.25 0.55 0.01 0.37 0.57 0.01 0.26 0.55
Right proposed 0.02 0.03 0.42 0.01 0.02 0.49 0.02 0.03 0.40
energy 0.00 0.00 0.22 0.01 0.00 0.25 0.01 0.00 0.22
@@ 6.2 Age-at-death distributions
The study of human mortality as quantified by age-at-death distributions has attracted much interest over the past decades. The analysis of age-at-death distributions has led to deeper insights into mortality patterns, lifespan variability and the factors influencing survival across different populations and time periods. The Human Mortality Database (www.mortality.org) provides annual age-at-death tables for 38 countries, with death rates recorded for each country across ages ranging from 0 to 110. By applying basic smoothing techniques to the yearly life tables available in this database, one may obtain smooth age-at-death density distributions indexed by both country and calendar year. We focus on 33 countries with available data spanning the years 1983 to 2018 and employ the conditional framework detailed in Section 4, where the data are considered to be triplets $(X,Y,Z)$. Here $X$ stands for the male age-at-death distribution for a given country at calendar year $Z$, and $Y$ the corresponding age-at-death distribution for females for the same country and year. We use calendar time as covariate $Z$, an inherently continuous variable, although it is initially observed only at discrete time points, taken to be the midpoints of calendar years. With 35 consecutive years in the dataset and mortality rates exhibiting smooth temporal variation, this observation scheme allows for the effective application of local smoothing methods in the pre-processing. We employ the Wasserstein-2 metric (16) in the space of one-dimensional age-at-death distributions, as it is particularly sensitive to horizontal shifts in distributions that are of special interest.
@@
We applied the proposed conditional profile association $\Delta_{XY}(z)$ in (12) for this analysis, aiming to shed light on the changing nature of the relationship between male and female age-at-death distributions across different calendar years. The conditional profile association serves to quantify the degree of dependence between the two distributions as a function of calendar time $z$, where a smaller value of $\Delta_{XY}(z)$ indicates that male and female death rate distributions are relatively less dependent, and vice versa.
@@
We divided the 33 countries into two groups: Eastern European countries (Bulgaria, Belarus, Czechia, Estonia, Hungary, Lithuania, Latvia, Poland, Slovakia, and Slovenia) and Western/Asian countries (Australia, Austria, Belgium, Canada, Switzerland, Denmark, Spain, Finland, France, England and Wales, Greece, Ireland, Iceland, Italy, Japan, Luxembourg, the Netherlands, Norway, New Zealand, Portugal, Sweden, Taiwan and the United States). The results are in Figure 5, which illustrates the estimated conditional profile association as a function of calendar year for all countries combined and separately for Eastern European and Western/Asian countries.
@@
The left panel, representing all countries, exhibits a generally decreasing trend over time, indicating a gradual reduction in the association between male and female age-at-death distributions over the 25 years covered in this analysis. The middle panel, which focuses on Eastern European countries only, shows a noticeable peak in the early 1990s, suggesting a period of heightened dependence between male and female mortality distributions, potentially due to major socio-political and economic changes in the region during that time that had a strong effect on both female and male age-at-death distributions. After this peak, the association declines but remains relatively stable. In contrast, the right panel, which represents Western/Asian countries, shows a consistently lower conditional profile association compared to both all countries combined and Eastern European countries. The association in Western countries also follows a steady downward trend, suggesting an increasing independence between male and female mortality distributions over time. This lower and declining association seems to indicate an increasing decoupling between male and female mortality over time that is more expressed for Western/Asian countries. The clear disparity between Western and Eastern European countries highlights the varying demographic, political and health-related forces that influence the relationship between male and female mortality.
To complement the observed decreasing trend in conditional profile associations between male and female mortality, we conducted the proposed conditional independence test described in Section 4. The resulting $p$-values are consistently below 0.01 for both Eastern European and Western/Asian countries, providing strong statistical evidence against conditional independence over the time period considered. These findings are expected and confirm that mortality rates for males and females are overall dependent, as expected.
@@ 7 Discussion
In this paper, we introduce the profile association, a measure designed to quantify associations between random objects. The proposed association is based on distance profiles, which serve as bridges connecting the distributions of random objects, the classical Hoeffding’s D statistic, and the proposed association measure. Utilizing this framework, we develop an independence testing procedure for random objects, demonstrating superior performance compared to existing methods in numerical studies. We further extend the profile association and the associated testing framework to conditional settings, in which random objects are paired with Euclidean predictors. This extension provides a useful tool for exploring relationships between complex object responses and Euclidean covariates, with immediate practical relevance. Beyond independence testing, the profile association framework has potential for broader statistical applications, including causal inference.
Causal inference has emerged as an important field in statistics, with significant applications across various scientific disciplines. One key challenge in causal inference is detecting confounding contributions that can lead to spurious associations arising from the influence of a confounder on two outcomes of interest. A simple conceptual tool for understanding confounding for Euclidean data is the classical law of total covariance,
@@
which decomposes the conditional covariance between random variables $X$ and $Y$ in the presence of a conditioning variable $Z$. Here $\text{Cov}(\mathbb{E}[X\mid Z],\mathbb{E}[Y\mid Z])$ is the contribution to the covariance explained by $Z$, and thus possibly due to a confounding effect of $Z$, whereas $\mathbb{E}[\text{Cov}(X,Y\mid Z)]$ is the part that is not explained by $Z$ in the context of a linear framework. A question for future research is whether by utilizing the proposed conditional profile association one can extend such a separation of the effect that can be attributed to $Z$ and the remaining association that cannot be explained by $Z$ to scenarios where $X$ and $Y$ are random objects and a linear framework does not apply, due to the lack of vector operations in non-Euclidean spaces.
Considering the mortality example discussed in Section 6.2, where $X_{i}$ and $Y_{i}$ are female and male age-at-death distributions, while $Z_{i}$ is a Euclidean covariate, a preliminary attempt to shed some light on this is the following three-step procedure: first compute the unconditional profile association $\Delta(X,Y)$, using all available data pairs $\{(X_{i},Y_{i})\}_{i}$; second, replace the conditional expectations $\mathbb{E}[X\mid Z]$, $\mathbb{E}[Y\mid Z]$ in the law of total covariance by Fréchet regressions (Petersen and Müller, 2019) to obtain conditional Fréchet means $\mathbb{E}_{\oplus}[X\mid Z]$ and $\mathbb{E}_{\oplus}[Y\mid Z]$ using data $\{(X_{i},Z_{i})\}_{i}$ and $\{(Y_{i},Z_{i})\}_{i}$ respectively; third, compute $\Delta(\mathbb{E}_{\oplus}[X\mid Z],\mathbb{E}_{\oplus}[Y\mid Z])$ by plugging in the results from the Fréchet regression step.
Applying these considerations for the human mortality data for all countries combined yields $30\Delta(X,Y)=0.36$ for the estimate of the overall profile association that ignores calendar year, but $30\Delta(\mathbb{E}_{\oplus}[X\mid Z],\mathbb{E}_{\oplus}[Y\mid Z])=0.97$ for the estimate of the part of the association explained by $Z$. Here we introduce the multiplier $30$ to normalize the association values so they lie in the interval $[0,1]$. Thus when calendar year is ignored, the observed profile association between male and female age-at-death distributions is much weaker than the value obtained when one first uses the calendar year as predictor. This discussion is not meant to be conclusive but rather to showcase the potential for future statistical research on theory and applications of conditional profile association for random objects.